Sonic the hedgehog is pushing a 7 ft long cart at 350 m/s. Midway through his run, he starts daydreaming about chili dogs and runs into a wall. Assume entire length of the cart is a "crumple zone."

1. Determine the acceleration applied to Sonic as he comes rest, assuming constant a.

2. Is the crash survivable (assume g limit is same as for humans, around 50 grams)
Answer #1

d = 2.13 m

a)

using 3rd equation of motion

a = v^2/ ( 2d)

a = 350^2/(2 x 2.13) = 28755.87 m/s^2

b)

a = 28755.87/9.8 = 2934.27 g

So, it is unsafe

Comment in case any doubt please rate my answer